OpenStudy (anonymous):

Need Help . ... Calculate centripetal acceleration & centripetal force on a man whose mass is 80kg who is resting on the ground at the Equator.

OpenStudy (anonymous):

as we know, Centripetal acceleration = V^2/r and Centripetal force = mv^2/r

OpenStudy (anonymous):

value of r will be is 6.4^*10^6 m in this case

OpenStudy (anonymous):

how to find the V ... ?

OpenStudy (anonymous):

how is that we arent flooting? how is that we are held by earth giving us centripoetal accleration?

OpenStudy (anonymous):

floating*

OpenStudy (anonymous):

dont know

OpenStudy (anonymous):

haerd of apple falling on newton because of GRAVITY?

OpenStudy (anonymous):

lol.. yeah..

OpenStudy (anonymous):

so? that garvitationa force provides centripetal acceleration

OpenStudy (anonymous):

hey how can u equate accleration(g) with velocity?

OpenStudy (anonymous):

gravitational force =GMm/r^2

OpenStudy (anonymous):

after that ?

OpenStudy (anonymous):

equate GMm/r^2=mv^2/r

OpenStudy (anonymous):

u understood why i equated both?

OpenStudy (anonymous):

No, I didnt understand can you explain further please

OpenStudy (anonymous):

in physics, when u say a body is accelerating in nature without u giving any accelartion, then that accleartion should have been provided by some external agency here in this case there is gravitational force acting between the earth and man which provides the centripetal force required for him to stay on earth *IMPORTANT: you shud always equate same quantities like accleration with acceleration velocity with velocity force with force what we are doing above is equating force with force

OpenStudy (anonymous):

ok, then what will er get after equating them. ?

OpenStudy (anonymous):

we*

OpenStudy (anonymous):

u get value of centripetal force directly=GMm/r^2 then divide it my massm to get value for centripeta acceleration as force=mass*accelaertaion

OpenStudy (anonymous):

Sorry, I am still not getting the correct answer, can you solve for that ?

OpenStudy (anonymous):

117373650919.5408*10^7 is centripetal acceleration or 1.173*10^18

OpenStudy (anonymous):

but the correct answer i'm provided is 0.0338m/s^2 how did you get that ?

OpenStudy (anonymous):

GMm/r^2 put everything in s.i units then the answer wat u wrote is acceleration to get is divide by mass(80) G=6.011*10^-11 M=6*10^24 R=6.4*10^6 m=80

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

Why I'm not getting you answer using my calculator :D ?

OpenStudy (anonymous):

your*

OpenStudy (anonymous):

u understood the concept right ? do it slowly.......MY ANSWER MAY NOT BE RIGHT!

OpenStudy (anonymous):

ok , will check it later. now gotta go for class .) and you please check your answer too, if you're free..

OpenStudy (anonymous):

ok